40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 271 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  3235   Fri Jul 16 13:05:48 2010 Kyung-haUpdateSUSLate update for 7/13 Tue (Tip Tilts)

[Jenne & Kyung-ha]

We suspended the mirror to one of the main frame with the ECD backplane we finished before. The hard task was to find the right balance for the mirror so that 1) it won't be tilted and 2) it'll be in the right position for the ECD backplanes so that the magnets attached to the mirror holder would be in the very center of each ECD holes. We used optical lever laser (red He/Ne) to check the balance of the mirror. We tried to use the jig for the mirror holder clamps but because of the size difference, we couldn't use it at all. (Since the magnets are very heavy, we thought the wire being not perfectly centered might work better. However, the jig dimension was way too different that the wire ended up in the middle of one of the holes.) Since there was no other clever way to attach the wire in the right position, we just tried to be as center/accurate as possible. After attaching wire to that mirror holder clamps, we hanged it to the frame. Again, we couldn't find any other accurate way to find the center so we held the wire and tried to adjust the mirror height as accurate as possible so that it can be in the right position in respect to ECD backplane and not be tilted at the same time. However, when we hanged the mirror, it was still tilted.. So we adjusted the mirror tilt using the mirror holder clamps. Since the holes on the clamps were ellipse shapes, we could adjust the position of the clamps a little bit. When we adjust the clamps, we started to tighten the screws when the mirror is NOT in the perfect position since the tightening up part changes the mirror angle anyways. Luckily, when we tightened up the last screw, the mirror was in the perfect position! After that, we poked the mirror several times to make sure that it comes back to the same place.

Amazingly, we could finish this whole hanging/adjusting process in about 30 mins! :D (Jan said it's because of his amazing moral support. :P Maybe he'll be there to support us everytime we work on the mirrors?)

  3234   Fri Jul 16 12:36:00 2010 Katharine, SharmilaUpdateeloglevitation

After last night's challenge (or inspiration), we levitated our magnet this morning.  Since the nice Olympus camera is not currently in the 40m, we had to use my less stellar camera, but despite the poor video quality you can still see the magnet returning to its stable equilibrium position.  Once we recover the better camera, we will post new videos.  Also, we haven't yet figured out how to put videos in line in the elog entry, so here are the youtube links:

 

levitation 1

levitation 2

 

We adjusted the gain on coil 1 so that the resistance from the pots was 57.1k (maximum gain of 101.2,).

currents from power supply, pre-levitation: 0.08 A and 0.34 A

post levitation: 0.08 A and 0.11 A


note: we're not sure why changing the gain on coil 3 changes the current through the power supply, so we'd like to investigate that next.

Attachment 1: CIMG0649.AVI
  3233   Thu Jul 15 23:51:47 2010 Mr. MaricHowToSUSLevitate me if you can

You guys must work harder.

mag_lev.jpg


  3232   Thu Jul 15 19:27:04 2010 ranaSummaryPSLRC trend after the insulation removal

Untitled.png

As you can see, there was not much (if any) worsening of the laser frequency fluctuation from removing the RefCav insulation. The plots below are zooomed in:

a.pngb.png

I have used the same peak-to-peak scale so that its easy to compare the fluctuations before (LEFT) and after (RIGHT).

As you can clearly see, the laser frequency moves just as much now (the SLOW_DC) as it did before when it had the insulation. Only now the apparent (i.e. fake) RC temperature fluctuations are much larger. So this sensor is fairly useless as configured.

  3231   Thu Jul 15 19:13:03 2010 ranaUpdateIOOMC_F check

Sometimes I like to plot the spectrum of MC_F. Its a good diagnosis of whether something is wrong.

The red trace is noisier than the blue reference. What is the cause of this?

Attachment 1: a.png
a.png
  3230   Thu Jul 15 16:57:31 2010 josephb, kiwamuUpdateCDSThe temporarily named c1scx machine getting ready to connect to IO chassis

This is the machine which will be the new x end front end machine.  Its IP is 192.168.113.86. 

We changed the root and controls passwords to the usual.  We have modified the controls user group to be 1001, by using "usermod -u 1001 controls" (we had to use the non-rtl kernel to get that command to work).

We changed /etc/fstab to point to /cvs/cds on Linux rather than some downs machine.  We added a link to /cvs/cds/rtcds in the local /opt directory.

We modified the /etc/rc.d/rc.local file to no longer run /opt/open-mx/sbin/omx_init start, /cvs/cds/geo/target/fb/mx_stream -d scipe12:0 om1, and /cvs/cds/geo/target/fb/mx_stream -d scipe12:0 -e2 -r2 om2.  We modified the /usr/bin/setup_shmem.rtl to run only c1x00 c1scx and c1spx.

I also commented out a line0 "/bin/rm -f /rtl*"

 

  3229   Thu Jul 15 16:16:51 2010 AlbertoConfigurationLSCPRC and SRC length adjustements

Quote:

Tell me whether it is correct or not. Otherwise I won't be able to sleep tonight.

Quote:

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.612 m

 Sorry. I was in a rush to go to the LIGO "all hands" meetings when I posted that elog entry, that I forgot a zero in the SRC length value. The correct values are:

dl(PRC) = -0.0266 m; dl(SRC) = 0.0612 m

The cavity absolute lengths are then:

L(PRC) = 0.5/2/f1*c - 0.0266 = 6.7466 m

L(SRC) = c/f2 + 0.0612 = 5.4798 m

where c is the speed of light; f1 = 11065399 Hz; f2 = 55326995 Hz

  3228   Thu Jul 15 15:57:10 2010 KojiConfigurationLSCPRC and SRC length adjustements

Tell me whether it is correct or not. Otherwise I won't be able to sleep tonight.

Quote:

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.612 m

  3227   Thu Jul 15 12:21:08 2010 AlbertoConfigurationLSCPRC and SRC length adjustements

Lately I've been trying to calculate the corrections to the recycling cavity lengths that would compensate for the phase that the sidebands will pick up from the arms in the upgraded interferometer.

To do that calculation , I tried two quite different ways, although equivalent in principle. They both use the optickle model of the 40m, but the calculation is made differently.

In the first way, I looked directly at the phases of the field: phase of [input field] / [reflected field], phase of [input field at PRM] / [transmitted field at SRM].

In the second way I looked at the demodulation phases of the LSC signals.

The first way is much simpler, especially from a computational point of view. It is the first I tried several weeks ago, but then I had abandoned because back then I thought it wasn't the correct way.

Anyway, both ways gave me the same results for the PRC length.
For the SRC length, the first way has given me a clear outcome. On the other hand, the second way has produced a less clear result.

According to these results, these would be the proposed adjustements to the cavity lengths:
dl(PRC) = -0.0266 m; dl(SRC) = 0.0612 m

I) 1st Way
a) case of arms ideal length (33.86 m)

sidebandPhaseRotation_73433447384.png sidebandPhaseRotation_73433447590.png


b) case arm length = 38.40 m

PRC

 sidebandPhaseRotation_73433447960.png zoom -> sidebandPhaseRotation_73433448864.png

SRC

sidebandPhaseRotation_73433445372.png zoom -> sidebandPhaseRotation_73433449354.png


II) 2nd Way
a) case of arms ideal length (33.86 m)

 demodPhaseVsRecCavLengths_73432637363.png


b) case arm length = 38.40 m

demodPhaseVsRecCavLengths_73433237349.png

  3226   Thu Jul 15 11:58:50 2010 josephbUpdateComputersAdded channel to ADCU_PEM (C0DCU1)

I modified the C1ADCU_PEM.ini file in /cvs/cds/caltech/chans/daq/ (after making a backup), and added a temporary channel called C1:PEM-TEMP_9, the 9 corresponding to the labeled 9 channel on the front of the BNC breakout in the 1Y7 rack.  The chnnum it was set to is 15008 (it was commented out and called C1:PEM-PETER_FE).  I also set the data rate to 2048.

I then did telnet fb40m 8087, and shutdown, and also hit the blue reconfig button on the DAQ status screen for the C0DCU1 machine.  The framebuilder came back up.  I confirmed the temporary channel, as well as the Guralp channels were still working from C0DCU1.

We have strung a cable in the cable trays from the SP table to the 1Y7 rack, which has been labeled as "Phasecam PD".  This will be used to record the output of an additional photodiode.

 

  3225   Wed Jul 14 20:15:04 2010 DmassBureaucracyCamerasIR Olympus

I borrowed the Olympus and forgot to leave a note - I should have it for at most a day. dmassey@ligo if you need it urgently

  3224   Wed Jul 14 19:36:17 2010 GopalUpdateOptic StacksExperimental Confirmation of COMSOL Analysis

I experimentally determined the spring constant of a single Vitol spring in order to obtain a rough estimate for the natural frequency of single-stack oscillation.

The procedure basically involved stacking metal bars of known mass onto the Vitol and using a caliper to measure deviations from the equilibrium length.

The plot below shows that, for small compressions, the response is linear to an R-squared of 0.98.

Untitled.png

The experimental spring constant came out to be about 270 lb/ft, or 3900 N/m.

Previous documents have listed that the top stack takes on a load of approximately 43 kg. per individual spring. A bit of calculation yields an experimental resonant frequency of 9.5 Hz.

Compared with the theoretical COMSOL first harmonic of about 7.5 Hz, there is a reasonable amount of error. Of course, I used this calculation as a simple ballpark estimate: errors from misplacement onto the Viton were minimized through use of a level, but were still inevitable on the mm scale. Since the two methods yield answers with the same order of magnitude, we are ready to move forward and build the remaining layers of the stack.

  3223   Wed Jul 14 19:15:26 2010 GopalSummaryOptic StacksREVISION: Eigenfrequency Analysis of Single Stack Complete

My previous eigenfrequency analysis was incorrect by two orders of magnitude due to the misuse of Young's Modulus information for Viton. After editing this parameter (as documented on 7/14 19:00), the eigenmodes became much more reasonable. I also discovered the Deformation option under the Surface Plotting Options, which makes the eigenmodes of the single stack much more apparant.

Attached are pictures of the first four eigenmodes:

First Eigenmode: y-translational, 7.49 Hz

Eigenfrequency_1_Stack4.png

Second Eigenmode: x-translational, 7.55 Hz

Eigenfrequency_2_Stack4.png

Third Eigenmode: z-rotational, 8.63 Hz

Eigenfrequency_3_Stack4.png

Fourth Eigenmode: z-translational, 18.26 Hz

Eigenfrequency_4_Stack4.png

 

Attachment 2: Eigenfrequency_2_Stack4.png
Eigenfrequency_2_Stack4.png
  3222   Wed Jul 14 19:00:56 2010 GopalConfigurationOptic StacksVitol Material Properties, REVISED

Quote:

Viton flouroelastomers have somewhat variable material properties. The following parameters are being used for eigenfrequency analysis.

Young's Modulus: 72,500-87,000 psi (cite: http://www.row-inc.com/pfa.html) *Accurate for PFAs

Poisson's Ratio: 0.48-0.50 (cite: http://www.engineeringtoolbox.com/poissons-ratio-d_1224.html) *Accurate for rubber

Density: 1800 kg/m^3 (cite: http://physics.nist.gov/cgi-bin/Star/compos.pl?matno=275)

 The Young's Modulus for PFA turned out to be orders of magnitude greater than for Viton. The revised values produced much more accurate results:

Young's Modulus for Viton-75: 1950 psi or 6.6 MPa

(Courtesy Row, Inc. and Steve Vass)

This website contains other details: http://www.row-inc.com/viton.html

  3221   Wed Jul 14 18:09:50 2010 josephb, razibUpdatePhase CameraSome cleanup behind 1Y2 rack of phasecamera electronics

We made an attempt at cleaning up the phase camera setup electronics.

We have moved a portion of the electronics onto the SP table (specifically the mixer, splitters, amplifiers, and associated power).  We put away a large number of cables which were unneeded, both BNC and power cables. The Innolight Mephisto power supply and one signal generator are still behind 1Y2 on top of a non-functioning VME crate.  The second VME crate was put along the south arm where two other VME crates already were.  We placed a fair number of BNC cables and power cords back on their cable racks or approriate storage space, so the rats nests of cables has been reduced.

We moved one power strip from plugging in beyind 1Y1, to the far side of the SP table (closer to the 1Y3 rack), and also found and plugged in another power strip (also on the far side of the SP table) and placed this underneath the SP table to be able to power the signal generator and Innolight Mephisto laser (its not plugged in currently, but we'd like to do so next week).

 

  3220   Wed Jul 14 16:39:06 2010 JenneUpdateDAQDebugging Guralp and reboots

[Joe, Jenne]

Joe got on the phone with Alex, and Alex's magic Alex intuition told him to ask about the RFM switch.  The C0DAQ_CTRL's overload light was orangeAlex suggested hitting the reset button on that RFM switch, which we did. That fixed everything -> c0dcu1 came back, as did the frame builder.  Rana had pointed out earlier that we could have brought back all of the other front ends, and enabled the damping of the optics even though the FB was still down.  It's okay to leave the front ends & watchdogs on, and just reboot the FB, AWG, and DAQ_CTRL computers if that is necessary.

Anyhow, once the FB was back online, we got around to bringing back all of the front ends (as usual, except for the ones which are unplugged because they're in the middle of being upgraded).  Everything is back online now.

After all of this craziness, all of the Guralp channels are working happily again. It is still unknown why they starting being digital zero, but they're back again. Maybe I should have rebooted the frame builder in addition to c0dcu1 last night?

 

Quote:

This is regards to zero signal being reported by the channels C1:PEM-SEIS_GUR1_X, C1:PEM-SEIS_GUR1_Y, and C1:PEM-SEIS_GUR1_Z.

I briefly swapped Guralp 1 EW and Guralp 2 EW to confirm to myself that it was not on the gurlap end (although the fact that its digital zero is highly indicative a digital realm problem).  I then unplugged the 17-32, and then 1-16 channel connections to the 110B.  I saw floating noise on the GUR2 channels, but still digital zero on the GUR1 channels, which means its not the BNC break out box.

There was a spare 110B, unconnected in the crate, so to do a quick test of the 110B, I turned off the crate and swapped the 110Bs, after copying the switch configuration of the first 110B to the second one.  The original 110B was labeled ADC 1, while the second 110B was labeled ADC 0.  The switches were identical except for the ones closest to the Dsub connectors on the front.  All those switches in that set were to the right, when looking down at the switches and the Dsub connectors pointing towards yourself.

Unfortunately, the c0duc1 never seemed to come up with the new 110B (ADC 0).  So we put the original 110B back.  And turned the crate back on. 

The fb then didn't seem to come back quite right.  We tried rebooting fb40m it, but its still red with status 1.  c0daqctrl is green, but c0dcu1 is red, although I'm not positive if thats due to fb40m being in a strange state.  Jenne tried a telnet in to port 8087 and shutdown, but that didn't seem to help.  At this point, we're going to contact Alex when he gets in around 12:30.

 

 

  3219   Wed Jul 14 13:03:04 2010 Gopal UpdateWIKI-40M Update7.8.10 - 7.14.10 Weekly Update

Summary of this Week's Activities:

Wed. 7/7: COMSOL Busbar tutorials; began stack design; began base; Viton rubber research

Thurs. 7/8: Completed Viton rubber research; updated materials; finished designing the base layer

Fri. 7/9: Research model coupling papers; extensive eLog entry about base design and troubleshooting

Sun. 7/11: Played around with Busbar to find first eigenfrequency; continued crashing COMSOL

Mon. 7/12: Intrusions in COMSOL eLog tutorial entry; research eigenfrequency analysis; successfully got first eigenmode of rectangular bar

Tues. 7/13: Updated Poisson ratio of Viton and subsequently succeeded in running eigenfrequency tests on base stack layer. Systematic Perturbation Tests were documented in the most recent elog entry. Discussed results with Rana and decided this didn't make sense. Analytical study required.

Wed. 7/14: Went over to machine shop to experimentally extrapolate spring constant of Viton. Calculations to be done in the afternoon.

  3218   Wed Jul 14 12:31:11 2010 nancyUpdateGeneralWeekly Update

Summary of this week's work Wednesday - Aligned the mode cleaner with Koji, and then measured the beam characteristics at MC2 end. Koji then taught me how to read the WFS signals Thursday - wrote a script to measure the signals and calculated the coefficients relating mirror movement and DC signals of WFS. To know the possibility of the control, found SVD of the coeff matrix, and condition number. Friday - Set up the measurement of QPD linear response using a laser outside the cavity. Took data. Monday - did the calculations and plotting for the above experiment. Then played around with the MEDM screens , and also tried to see what happens to the Power Spectrum of WFS signals by changing the coefficients in teh matrix. (failed) Tuesday - played around with WFS, tried seeing what it does when switched on at different points, and also what it does when I disturb the system while WFS has kept it locked.

Today - had switched off the WFS sensors yesterday night after locking the MC as wanted to know that how does MC behave when no WFS gain is applied. I checked in the morning, the MC was locked all night. I am now proceding with my calculations for the sensing matrix
  3217   Wed Jul 14 12:12:03 2010 RazibSummaryPhase CameraWeekly update

This week I was mainly interested in investigating the noise source at the phase camera. So having this issue in mind, my activities are the following:

1. I worked on producing multiple beat signal (1Hz and 5Hz). Elog entry.

2. I altered the setup so that instead of triggering the camera from the signal generator, we are now triggering it from the beat signal from the reference beam and sideband.

3. I made the nice little aluminium table for all the amplifiers, mixer and splitters to sit at one place instead of floating around.

4. I talked with Aidan and Joe and verified my calculation and extended it to further investigation of the noise source in the setup.

 

Plan for the upcoming week:

1. Measure and calibrate the camera w.r.t the power incident on it.

2. Investigate the noise source.

  3216   Wed Jul 14 11:54:33 2010 josephbUpdateDAQDebugging Guralp and reboots

This is regards to zero signal being reported by the channels C1:PEM-SEIS_GUR1_X, C1:PEM-SEIS_GUR1_Y, and C1:PEM-SEIS_GUR1_Z.

I briefly swapped Guralp 1 EW and Guralp 2 EW to confirm to myself that it was not on the gurlap end (although the fact that its digital zero is highly indicative a digital realm problem).  I then unplugged the 17-32, and then 1-16 channel connections to the 110B.  I saw floating noise on the GUR2 channels, but still digital zero on the GUR1 channels, which means its not the BNC break out box.

There was a spare 110B, unconnected in the crate, so to do a quick test of the 110B, I turned off the crate and swapped the 110Bs, after copying the switch configuration of the first 110B to the second one.  The original 110B was labeled ADC 1, while the second 110B was labeled ADC 0.  The switches were identical except for the ones closest to the Dsub connectors on the front.  All those switches in that set were to the right, when looking down at the switches and the Dsub connectors pointing towards yourself.

Unfortunately, the c0duc1 never seemed to come up with the new 110B (ADC 0).  So we put the original 110B back.  And turned the crate back on. 

The fb then didn't seem to come back quite right.  We tried rebooting fb40m it, but its still red with status 1.  c0daqctrl is green, but c0dcu1 is red, although I'm not positive if thats due to fb40m being in a strange state.  Jenne tried a telnet in to port 8087 and shutdown, but that didn't seem to help.  At this point, we're going to contact Alex when he gets in around 12:30.

 

  3215   Wed Jul 14 11:51:48 2010 RazibUpdatePhase CameraWork near 1Y2 yesterday

Quote:

Razib and I were attempting to get the output of a photodiode (PD55A in this case) recorded, so that we could independently measure the slow (~1-10 Hz) fluctuations of the light incident on the camera.  This would then allow us to subtract those fluctuations out, letting us get at the camera noise in the case with signal present (as opposed to just a dark noise measurement when we look at the noise with no signal present).

Originally I was thinking of using one empty patch panel BNCs used for PEM channels down by the 1Y7 rack and go through a 110B, although Alberto pointed out he had recently removed some monitoring equipment, which watched the amplitude modulation at various frequencies of the RF distribution (i.e. 33 MHz, etc).  This equipment output a DC voltage proportional to the amplitude of the RF signals.  The associated channel names were C1:IOO-RFAMPD_33MHZ, C1:IOO-RFAMPD_33MHZ_CAL, C1:IOO-RFAMPD_133MHZ, etc.  These are slow channels, so I presume they enter in via the slow computers, probably via pentek (I didn't check that, although in hindsight I probably should have taken the time to find exactly where they enter the system).  The connections them selves were a set of BNCs on the south side, half way up the 1Y2 rack.

We simply chose one, the 33 MHz channel in this case, and connected.  At around this time, the MC did become unlocked, although it looked like it was due to the MC2 watchdog tripping.  The initial theory was we had bumped the Mode Cleaner while looking around for some BNC cables, although from what Rana had to do last night, it probably was the connection.  We were able to restore the watchdog and confirm that the optic started to settle down again.  Unfortunately, I had to leave about 5 minutes later, and didn't do as thorough an investigation as was warranted.

 Before I left, I disconnected the PD55, so the 33 MHz channel wasn't physically connected to anything!!! Only one end of the wire was connected to the rack while the other was free...

So it wasn't the PD connection that is responsible for MC tripping at the later time...

  3214   Wed Jul 14 11:32:36 2010 Sharmila, KatharineUpdateWIKI-40M Update 

Weekly update:

We correctly connected our circuit to power source to verify that it was functional and that our LED in the shadow sensor turned on.  It did, but we also noticed a funky signal from the LED driver.  We continued to attempt 1x1 levitation, but determined that the temporary flag we were using out of convenience (a long, thin cylindrical magnet) was weakly attracted to residually magnetized OSEM components.  We then switched to an aluminum screw as our flag.

We resoldered and applied heat shrink to the wires connecting our coil to the BNC terminal, since they were falling apart.

We sat down with Rana and removed circuit components in the LED drive part by part to determine what was tripping up the circuit.  We determined a rogue capacitor to be at fault and removed it from the circuit.

We used a spectrum analyzer to measure the frequency response of our circuit (see details in last elog).  We are currently making a Simulink block diagram so we can check the stability of our setup, but are temporarily set back because our plotted calculation of the transfer function clearly doesn't match the measured one.

  3213   Wed Jul 14 10:00:14 2010 josephbUpdatePhase CameraWork near 1Y2 yesterday

Razib and I were attempting to get the output of a photodiode (PD55A in this case) recorded, so that we could independently measure the slow (~1-10 Hz) fluctuations of the light incident on the camera.  This would then allow us to subtract those fluctuations out, letting us get at the camera noise in the case with signal present (as opposed to just a dark noise measurement when we look at the noise with no signal present).

Originally I was thinking of using one empty patch panel BNCs used for PEM channels down by the 1Y7 rack and go through a 110B, although Alberto pointed out he had recently removed some monitoring equipment, which watched the amplitude modulation at various frequencies of the RF distribution (i.e. 33 MHz, etc).  This equipment output a DC voltage proportional to the amplitude of the RF signals.  The associated channel names were C1:IOO-RFAMPD_33MHZ, C1:IOO-RFAMPD_33MHZ_CAL, C1:IOO-RFAMPD_133MHZ, etc.  These are slow channels, so I presume they enter in via the slow computers, probably via pentek (I didn't check that, although in hindsight I probably should have taken the time to find exactly where they enter the system).  The connections them selves were a set of BNCs on the south side, half way up the 1Y2 rack.

We simply chose one, the 33 MHz channel in this case, and connected.  At around this time, the MC did become unlocked, although it looked like it was due to the MC2 watchdog tripping.  The initial theory was we had bumped the Mode Cleaner while looking around for some BNC cables, although from what Rana had to do last night, it probably was the connection.  We were able to restore the watchdog and confirm that the optic started to settle down again.  Unfortunately, I had to leave about 5 minutes later, and didn't do as thorough an investigation as was warranted.

  3212   Wed Jul 14 01:05:27 2010 Sharmila,KatharineSummaryelogMaglev

Yesterday we hooked up the Quadrant Maglev control to the power supply to test the components in the Input/Output part of the circuit.

The output from the buffer was an unexpected high noise signal which was caused by some circuit components.

Consequently these were replaced/removed after confirming the source of noise.

The following is a story of how it was done.


To test the components of input/output, we measured the output across TP_PD3(Test Point -Photo Diode 3).
We got a high noise signal with a frequency of several kHz.

We tested the values of various electronic components. The resistances R5 and R6 did not measure as mentioned(each had a value of 50 K in the schematic). The value of R6 was 10 K and we replaced R5 with a 10 K resistor. We still got the noise signal at 5.760 kHz with a Pk-Pk voltage of 2.6 V. The resistors in R-LED measured 1.5 K instead of the marked 2.2 K.

P7120278.JPG
We had three suspects in hand:

  • BUF634P : A buffer from the Sallen-Key filter to the LED.
  • C24 : A capacitor which is a part of the Sallen-Key filter.
  • C23 : A capacitor in the feedback circuit of the Sallen-Key filter.


BUF634P : The data sheet for the BUF634P instructed a short across the 1-4 terminals in the presence of capacitive load.  We followed this to overcome the effect(if any) of the extra-long BNC cables which we were using. The oscilloscope still waved 'Hi!' at a few kHz. We removed the buffer and also the feedback resistor R42 from the circuit, what we were testing boiled down to measuring the output of the Sallen-Key filter. The output still contained the funny yet properly periodic signal at a few kHz.      

.P7120284.JPG    


C24: Removing C24 did not do any good.

C23: As a final step C23 was removed. And ... We got a stable DC at 9.86 V(almost stable DC with a low noise at a few MHz). C24 and the buffer were replaced and output seemed fine. The output was a high frequency sine wave which was riding on a DC of 9.96 V.

 

P7120281.JPG


We rechecked if the LED was on and the infrared viewer gave a positive signal.



We went ahead obtaining the transfer function of the feedback control for which we used a spectrum analyzer.

The input for feedback system is a photo current whereas the spectrum analyzer tests the circuit with a voltage impulse.  Hence the voltage input from the spectrum analyzer needs to be converted into current of suitable amplitude(few microamps) for testing the spectrum analyzer.  Similarly the output which is a coil current needs to be changed to a voltage output through a load for feeding into the channel of the spectrum analyzer. We used a suitable resistance box with BNC receiving ends to do this. We obtained a plot for the transfer function which is shown below.

P7140292.JPG


Future plans:

- Check the calculated transfer functions with the plot of the spectrum analyzer

- Model the entire(OSEM, magnet, actuators etc.) system in Simulink and calculate the overall transfer function

- Stable levitation of the 1X1 system

  3211   Tue Jul 13 23:20:37 2010 JenneUpdatePEMGur1 has no signal?!?!?!?!

Quote:

[Sanjit, Jenne]

Sanjit discovered that the Gur1 channels are all digital 0.  We determined that this began on July8, 04:00 UTC (~9pm on the 7th?).

It's digital zero, so we suspect a software thing.  Just to check, we put a sine wave in, and didn't see anything.  Gur2 seems totally fine, and the sine wave input showed up nicely on dataviewer.  What's going on? Sabotage to prevent this paper from getting done?  Dmass trying to get his paper done before me???

Investigations are ongoing.... Joe claims it's not his fault, since his shenanigans near the PEM rack were on days before, and days after this, but not on the 7th.  

 I just rebooted c0dcu1, which didn't help anything.  Joe said he'd try to give me a hand tomorrow.

  3210   Tue Jul 13 21:04:49 2010 tara,ranaSummaryPSLTransfer function of FSS servo

 I measured FSS's open loop transfer function.

For FSS servo schematic, see D040105-B.  

4395A's source out is connected to Test point 2 on the patch panel.

Test Point 2 is enabled by FSS medm screen.

"A" channel is connected to In1, on the patch panel.

"R" channel is connected to In2, on the patch panel.

the plot shows signal from A/R.

Note that the magnitude has not been corrected for the impedance match yet.

So the real UGF will be different from the plot.

 

-------------------------

4395A setup

-------------------------

network analyzer mode

frequency span 1k - 10MHz

Intermediate frequency bandwidth 100Hz

Attenuator: 0 for both channels

Source out power: -30 dBm

sweep log frequency

------------------------------

medm screen setup

-----------------------------

TP2: enabled

Common gain -4.8 dB

Fast Gain 16 dB

Attachment 1: TF_FSS_ser.png
TF_FSS_ser.png
  3209   Tue Jul 13 19:26:47 2010 ranaUpdateIOOMC still broken

After whatever Joe/Alberto did this afternoon, the MC was not locking. Koji and I removed several of the cables in the side of the rack where they

were apparently working (I say apparently because there's no elog).

MC is now locking but the autolocker did not work at first - op340m was unable to access any channels from c1iool0. After several minutes, it mysteriously

started working - the startup.cmd yields errors seen on the terminal. I attach the screen dump/.

Attachment 1: a.txt.gz
  3208   Tue Jul 13 17:36:42 2010 nancyUpdateIOOWavefront Sensing Matrix Control

For yesterday - July 12th.

Yesterday, I tried understanding the MEDM and the Dataviewer screens for the WFS.

I then also decided to play around with the sensing matrix put into the WFS control system and see what happens.

I changed the sensing matrix to completely random values, and for some of the very bad values, it even lost lock :P (i wanted that to happen)

Then I put in some values near to what it already had, and saw things again.

I also put in the matrix values that I had obtained from my DC calculations, which after Rana's explanation, I understand was silly.

Later I put back the original values, but the MC lock didnot come back to what it was earlier. Probably my changing the values took it out of the linear region. THE MATRIX NOW HAS ITS OLD VALUES.

I was observing the POwer Spectrum of teh WFS signals after changing the matrix values, but it turned out to  be a flop, because  I had not removed the mean while measuring them.  I will do that again today, if we obtain the lock again (we suddenly lost MC lock badly some 20 minutes ago).

  3207   Tue Jul 13 14:59:04 2010 GopalUpdateOptic StacksEigenfrequency Analysis of Single Stack Complete

Via reconfiguration of Viton parameters (previously posted), I managed to debug the COMSOL run time errors and null pointer exceptions. Listed are the resultant eigenfrequencies obtained through structural analysis testing. For all tests, the bottom of the Viton springs are constrained from motion, and all other parts are free to oscillate. Notice that color variations signify displacement from the equilibrium position. Also note that different initial conditions produce different eigenmodes:

No initial displacement:

Eigenmode_Stack_4.png


0.01 m x-displacement:

Eigenmode_Stack_4_xdisp.png


0.01 m y-displacement:

Eigenmode_Stack_4_ydisp.png


 0.01 m z-displacement:

 Eigenmode_Stack_4_zdisp.png


Clearly, the plate has its first harmonic between 210-215 Hz, which is much greater than seismic noises (which never exceed the 10-Hz range). This suggests a highly attenuating transfer function. Since the remaining three plates have been designed to resonate similarly, it is likely that the entire stack system will also function very well.

Next steps:

1) Extend the eigenfrequency analysis to obtain a transfer function for the single-plate system

2) Expand the CAD model to include all four stack layers, and perhaps a base

 

  3206   Tue Jul 13 13:56:29 2010 GopalConfigurationOptic StacksVitol Material Properties

Viton flouroelastomers have somewhat variable material properties. The following parameters are being used for eigenfrequency analysis.

Young's Modulus: 72,500-87,000 psi (cite: http://www.row-inc.com/pfa.html) *Accurate for PFAs

Poisson's Ratio: 0.48-0.50 (cite: http://www.engineeringtoolbox.com/poissons-ratio-d_1224.html) *Accurate for rubber

Density: 1800 kg/m^3 (cite: http://physics.nist.gov/cgi-bin/Star/compos.pl?matno=275)

  3205   Tue Jul 13 11:51:19 2010 JenneUpdatePEMGur1 has no signal?!?!?!?!

[Sanjit, Jenne]

Sanjit discovered that the Gur1 channels are all digital 0.  We determined that this began on July8, 04:00 UTC (~9pm on the 7th?).

It's digital zero, so we suspect a software thing.  Just to check, we put a sine wave in, and didn't see anything.  Gur2 seems totally fine, and the sine wave input showed up nicely on dataviewer.  What's going on? Sabotage to prevent this paper from getting done?  Dmass trying to get his paper done before me???

Investigations are ongoing.... Joe claims it's not his fault, since his shenanigans near the PEM rack were on days before, and days after this, but not on the 7th.  

Attachment 1: Gur1_noSignal_13Jul2010.png
Gur1_noSignal_13Jul2010.png
  3204   Tue Jul 13 11:20:07 2010 DmassUpdateGreen LockingSHG on PSL table : optimum temeprature

 

 It seems like you might inherit an offset by using step (3) b/c of the temperature gradient between the crystal and the sensing point. Depending on how large this gradient is you could increase the linear coupling from temperature to intensity noise from zero to a significant number. Phase noise should not be effected.

SInce these things (ovens) are so low time constant, shouldn't we

  1. Lock to a temperature
  2. Let the oven equilibrate for however long - a few tau maybe - my oven has a time constant of 60 sec, don't know if this is fast or slow compared to that
  3. Measure P_532/P_1064
  4. Change the setpoint
  5. Go back to step 1
  3203   Tue Jul 13 11:00:29 2010 kiwamuUpdateGreen LockingSHG on PSL table : optimum temeprature

The optimum temperature for the doubling crystal on the PSL table was found to be 36.8 deg.

I scanned the temperature of the crystal from 44 deg to 29 deg, in order to find the optimum temperature where the frequency doubled power is maximized. 

 

(method) 

The method I performed is essentially the same as that Koji did before (see this entry).

(1) First of all, I enabled the PID control on the temperature controller TC200 and set the temperature to 44 deg.

(2) After it got 44 deg, I disabled the PID control.

(3) Due to the passive cooling of the oven, the temperature gradually and slowly decreased. So it automatically scans the temperature down to the room temperature.

(4) I recorded the power readout of the power meter: New Port 840 together with the temperature readout of TC200. The power meter was surely configured for 532 nm.

 

(result)

The measured data are shown in the attachment. 

The peak was found at T=36.8 deg where the power readout of  532 nm was 605 uW.

Compared with Koji's past data (see this entry), there are no big side lobes in this data. I am not sure about the reason, but the side lobes are not critical for our operation of the green locking.

 

 (to be done)

 Adjustment of the PID parameters

Attachment 1: power_temp.png
power_temp.png
  3202   Tue Jul 13 10:02:30 2010 steveUpdateMOPAlaser power is dropping slowly

I have just removed an other 400 cc of water from the chiller.  I have been doing this since the HTEMP started fluctuating.

The Neslab bath temp is 20.7C, control room temp 71F

 

Attachment 1: power100d.jpg
power100d.jpg
  3201   Mon Jul 12 22:01:13 2010 KojiUpdateSUSOne TT suspended. Still need fine alignment

Jenne and Koji

We tweaked the alignment of the TT mirror.

First we put a G&H mirror, but the mirror was misaligned and touching the ECD as the magnet was too heavy. We tried to move the wires towards the magnet by 1mm.
It was not enough but once we moved the clamps towards the magnet, we got the range to adjust the pitching back and forth.
We tried to align it by the feaher touch to the clamp, we could not get close to the precision of 10mrad as the final tightening of the clamp screws did change the alignment.

We will try to adjust the fine alignment tomorrow again.

The damping in pitch, yaw and longitudinal looks quite good. We will also try to characterize the damping of the suspension using a simple oplev setup.

Attachment 1: IMG_2634.jpg
IMG_2634.jpg
  3200   Mon Jul 12 21:26:02 2010 Katharine, SharmilaUpdateelogmaglev coils

The connection between our coil wires and BNC terminals was pretty awful (soldered wires broke off ) so we removed the old heat shrink and re-soldered the wires.  We then chose more appropriately sized heat shrinks (small one around each of the two soldered wires, a medium-sized shrink around the wires together, a large one covering the BNC terminal and the wire) and used the solder iron and heat gun to shrink them.

P7120276.JPG

 

 

  3199   Mon Jul 12 18:37:10 2010 GopalUpdateOptic StacksEigenfrequency Analysis of Simple Objects

Eigenfrequency analysis has been successfully completed in COMSOL on both a tutorial camshaft, as well as a homemade metal bar.

Upon increasing in complexity to the busbar, I once again began getting into run time errors and increased lag. It seems that this is due to undefined eigenvalues when solving the linear matrices. I tried many boundary values as well as initial conditions in case this was the issue, but it was not. There seems to be some sort of an internal inconsistency. This is no longer a matter of tweaking parameters.

Next steps:

1) Try using the same techniques on the actual mirror stacks to see if we get lucky.

2) In the likely case that this doesn't happen, continue the debugging process. If necessary, a good deal of time may need to be spent learning the COMSOL lower-level jargon.

  3198   Mon Jul 12 17:05:30 2010 nancyUpdateSUSCharacterisation of the QPD

Quote:

I and koji setup the measurement of the QPD response to the pitch and yaw displacements of the beam spot.

We did this using a 100mW 1064nm laser. Its power was attenuated to ~ 1.9mW, and the spot size at the QPD position was 6000-7000 um .

The QPD was put on a translation stage, using which, the center of teh QPD wrt the beam spot could be moved in pitch and yaw.

Following are the measurements :

 

 The old plots looked horrible, and so here is a new plot

plot.png

The slopes and other stats are

Pitch

Linear model Poly1:
     f(x) = p1*x + p2
Coefficients (with 95% confidence bounds):
       p1 =        8550  (7684, 9417)
       p2 =       -2148  (-2390, -1906)

Goodness of fit:
  SSE: 9944
  R-square: 0.9923
  Adjusted R-square: 0.9907
  RMSE: 44.59

Yaw

Linear model Poly1:
     f(x) = p1*x + p2
Coefficients (with 95% confidence bounds):
       p1 =       -8310  (-8958, -7662)
       p2 =        2084  (1916, 2252)

Goodness of fit:
  SSE: 6923
  R-square: 0.9954
  Adjusted R-square: 0.9945
  RMSE: 37.21

Attachment 1: plot.png
plot.png
  3197   Mon Jul 12 15:49:56 2010 nancyUpdateSUSCharacterisation of the QPD

I and koji setup the measurement of the QPD response to the pitch and yaw displacements of the beam spot.

We did this using a 100mW 1064nm laser. Its power was attenuated to ~ 1.9mW, and the spot size at the QPD position was 6000-7000 um .

The QPD was put on a translation stage, using which, the center of teh QPD wrt the beam spot could be moved in pitch and yaw.

Following are the measurements :

For yaw

:fullyaw.jpg

The slope of teh linear region is -8356 /inch

yaw_linear.jpg

 For pitch

fullpitch.jpg

The slope of the linear region in this is 9085/inch

 

pitch_linear.jpg

 

  3196   Mon Jul 12 14:22:36 2010 JenneSummaryPSLThings I did to the PSL today: Refcav, PMC, cameras, etc.

Quote:

I re-aligned the beam into the PMC. I got basically no improvement. So I instead changed the .LOW setting so that PMCTRANS would no longer go yellow and make the donkey sound.

I did the same for the MOPA's AMPMON because its decayed state is now nominal.

[Jenne, Chip]

The alarm was still going, because the LOLO setting was higher than the LOW, which is a little bit silly.  So we changed the .LOLO setting to 0.80 (the LOW was set to 0.82)

We also changed psl.db to reflect these values, so that they'll be in there the next time c1psl gets rebooted.

  3195   Mon Jul 12 13:16:53 2010 kiwamuUpdateGreen LockingPZT feedback at X end

The feedback signal going to the laser PZT at the X end station was measured in the daytime and the nighttime.

It's been measured while the laser frequency was locked to the arm cavity with the green light.

arm_day_night.png

 

The red curve was measured at 3pm of 8/July Friday. And the blue curve was measured at 12am of 9/July Saturday. 

As we can see on the plot, the peak-peak values are followers

              daytime:  ~ 4Vpp

       nighttime:  ~1.8Vpp

It is obvious that the arm cavity is louder in the daytime by a factor of about 2.

Note: the feedback signal is sent to the PZT only above 1Hz because the low frequency part is stabilized mostly by the crystal temperature (see this entry)

Quote:

 What we care about is the peak-peak value of the PZT feedback signal measured on a scope for ~30 seconds.

  3194   Mon Jul 12 12:16:50 2010 DmassHowToCOMSOL TipsIntrusions (Negative Extrusions)

 An entry on the 40m wiki page might serve you better, and be easier to sift through once all is said and done

  3193   Mon Jul 12 11:20:56 2010 Gopal HowToCOMSOL TipsIntrusions (Negative Extrusions)

For the sake of future users, I have decided to periodically add tips and tricks in using COMSOL that I have figured out, most probably after hours of circuitous efforts. They will always be listed under the new COMSOL Tips category.

Today's topic: Intrusions

COMSOL has a very user-friendly interface for taking objects from 2D to 3D using the "extrusion" feature. But suppose one wants to design an object which contains screw holes or some other indentation. I've found that creating "punctures" in COMSOL is either impossible or very complicated.

Instead, COMSOL encourages users to always "add" to the object. In other words, one must form the lowest level first, then build layers sequentially on top using new work plane and boolean difference operators. This will probably be a bit clearer with an example:

1) First, create the planar projection in a work plane:

Screen_shot_2010-07-12_at_10.51.22_AM.png

2) Extrude the first layer only in the regular fashion:

Screen_shot_2010-07-12_at_10.51.28_AM.png

 3) Add a new work plane which is offset in the z-direction to the deepest point of the intrusion.

Screen_shot_2010-07-12_at_10.52.08_AM.png

 4) Now, create the shape of the intrusion in this new work plane.

Screen_shot_2010-07-12_at_10.53.53_AM.png

5) Use the Boolean "Difference" to let COMSOL know that, on this plane, the object has a hole.

 Screen_shot_2010-07-12_at_10.54.36_AM.png

 6) Extrude once more from the second work plane to complete the intrusion.

Screen_shot_2010-07-12_at_10.55.36_AM.png

  3192   Mon Jul 12 10:23:51 2010 steveUpdatePEMair condition maintenance is today

The AC filters will be checked and/or replaced today. This means the AC will be off for sort periods of time. Temperature and particle count will be effected some what.

See 800 days plot

Attachment 1: pem2y.jpg
pem2y.jpg
  3191   Mon Jul 12 02:21:01 2010 KojiConfigurationASCResurrection of MC WFS

I have resurrected the MC WFS on Friday night.
I have uncommented the WFS part of the MC autolocker.
The WFS total gain was empirically set to 0.1 such that the loops have no instability.

The loops somewhat worked through the weekend although they seemed to have the drift of the operating points
in accordance with the WFS spot.

  3190   Sun Jul 11 20:11:48 2010 ranaSummaryPSLRC trend after the insulation removal

rc.png

 

  3189   Fri Jul 9 20:16:19 2010 ranaSummaryPSLThings I did to the PSL today: Refcav, PMC, cameras, etc.

I re-aligned the beam into the PMC. I got basically no improvement. So I instead changed the .LOW setting so that PMCTRANS would no longer go yellow and make the donkey sound.

I did the same for the MOPA's AMPMON because its decayed state is now nominal.

 

Steve and I removed the thermal insulation from around the reference cavity vacuum chamber. It wasn't really any good anyways.

Here are the denuded photos:

 

Steve and I are now planning to replace the foam with some good foam, but before that we will wrap the RC chamber with copper sheets like you would wrap a filet mignon with applewood bacon.

This should reduce the thermal gradients across the can. We will then mount the sensors directly to the copper sheet using thermal epoxy. We will also use copper to cover most of this hugely

oversized window flange - we only need a ~1" hole to get the 0.3 mm beam out of there.

 

My hope is that all of this will improve the temperature stability of this cavity. Right now the daily frequency fluctuations of the NPRO (locked to the RC) are ~100 MHz. This implies

that the cavity dT = (100 MHz) / (299792458 / 1064e-9) / (5e-7) = 1 deg.    That's sad....

 

I also changed the RC_REFL cam to manual gain from AGC. I cranked it to max gain so that we can see the REFL image better.

  3188   Fri Jul 9 12:25:25 2010 kiwamuUpdateGreen LockingSHG on PSL table

In order to increase the green power on the PSL table, I moved the position of the Second Harmonic Generation (SHG) crystal by ~5cm.

After this modification, the green power increased from 200 uW to 640 uW. This is sufficiently good.

     As I said in the past elog entry (# 3122), the power of the green beam generated at the PSL table should be about 650 uW.

I measured the green power by the Ophir power meter and found it was ~200 uW, which made me a little bit sad.

Then I performed the beam scan measurement to confirm if the crystal  was  located on the right place. And I found the postion was off from the optimum position by ~5cm.

So I slided the postion of the SHG oven to the right place and eventually the power got increased to 640 uW.

 

some notes: 


(power measurement)

        The outgoing beam from the SHG crystal is filtered by Y1-45S to eliminate 1064nm.

According to Mott's measurement Y1 mirrors are almost transparent for green beams (T~90%), but highly reflective for 1064nm (T~0.5%).

All the green power were measured after the Y1 mirror by the Ophir configured to 532nm, though, the measured power might be offseted by a leakage of 1064nm from the Y1 mirror.

I didn't take this effect into account.

 

 

(beam scanning and positioning of crystal)

          Here is the properties of the incident beam. These numbers are derived from the beam scan measurement.

w0h             = 52.6657      +/- 0.3445 um

w0v             = 52.4798      +/- 0.1289  um

z0h             = 0.574683         +/- 0.001937 m

z0v             = 0.574325         +/- 0.0007267 m

Where the suffixes "h" and "v" stand for "horizontal" and "vertical" respectively.

The distances are calibrated such that it starts from the lens postion.

Both waist size are already sufficiently good because the optimum conversion can be achieved when the waist size is about 50um ( see this entry)

The measured data and their fitting results are shown in attachement 1.

         According to my past calculation the center of the crystal should be apart from the beam waist by 6.8mm (see this entry)

So at first I put the oven exactly on the waist postion, and then I slided it by 6.8mm.

 

 

(to be done)

        I need to find an optimum temperature for the crystal in order to maximize the green power.

Previously the optimum temperature for the crystal was 38.4 deg. But after moving the position I found the optimum temperature is shifted down to around 37deg.

Attachment 1: PSL_doubling.png
PSL_doubling.png
  3187   Fri Jul 9 12:07:26 2010 RazibUpdatePhase CameraWeekly update

Here are some more details about the current phasecam setup. We are using a He-Ne laser setup

phase_cam_setup_09_08_10.jpg

A crude snap shot of the setup....

mod_setup_(copy)_annotated.jpg

 

We sent light through SM2 (Steering Mirror 2)  to BS1(Beam-Splitter 1) where the laser light is split into two parts, one going to the AOM and the other to the EOM. The EOM adds 40 MHz sidebands to the incoming carrier light after SM3, and the AOM shifts the frequency of the incident light on it to 40.000 005 MHz. The purpose for doing this juggling is to intentionally create a beat signal off the reference beam from the AOM with the sidebands added at the EOM. Note that, we are driving the AOM at 7dBm and the EOM at 13 dBm with 0 (nil) modulation. The two lights are combined at the BS2 and sent off through SM5 to the camera. The CMOS of the camera contains silicon based Micro MT9V022 chip which has a quantum efficiency of 50% at 633 nm. Thus we expected a fairly good response to this He-Ne setup from the camera. 

Using a trigger circuit, we triggered the camera at 20 Hz by sending a 20Hz sinusoidal signal to it. The trigger circuit converts this to a positive square waves. Then I roughly figured out the optimum exposure time for the camera before the DC levels saturates it by writing a code for taking a series of 25 images at different exposure time. I found that 500µs seems to be a reasonable exposure time. So, using this data, I took 20 consecutive images and sent them through a short Fourier Transform segment using Matlab to see the beat signal. Note that the DC component from these processing of the images have some fringe pattern which is due to the ND 2.5 filter that we were using to reduce the light power incident on the camera. The FT method also gave us the presence of the beat signal at the corresponding bin of the FT (for example: for 5Hz beat signal, I got the DC at bin 1 of the FT and 5Hz component at bin 6 as expected). Then I changed the AOM driving frequency to 40.000 001 MHz in order to see a 1 Hz beat signal. The results for both is in my previous post. 

Quote:

I have completed the following tasks:

1. Find a simplified calibration of the exposure time for the current He-Ne setup. Basically, I triggered the camera to take 20 snapshots with a 20 Hz driving signal at different exposure time beginning from 100 us (microsecond) upto 4000 us with an increment of 200 us.

    Result: The current power allows the camera to have an exposure time of ~500 us before the DC level begans to saturate.

2. Aidan and I did some alignment and connected the AOM and corrected the driving frequency of its PZT to 40 Mhz(which apparently was disconnected which in turn gets the credit of NO beat signal for me until Tuesday 07/06/2010 5:30 PST) .

    Result: I got the beat signal of 1 Hz and 5 Hz. Image follows (the colormap shows the power in arbitrary units).

3. Finished writing my Progress Report 1 .

DC_1Hz_beat_sig.jpgDC_5Hz_beat_sig.jpg

 

  3186   Fri Jul 9 11:41:58 2010 GopalSummaryOptic StacksTop Optic Layer Complete; Mechanical Tests Giving Problems

For the last week, I have been attempting to determine the mirror stack transfer function which relates mechanical mirror response to a given ground-motion drive. The idea is to model the stacks in COMSOL and ultimately apply mechanical tests for manual calculation.

 

Procuring the correct drawings to base my 3D models off of was no simple task. There are two binders full of a random assortment of drawings, and some of them are associated with the smaller chambers, while others are appropriate for the main chamber, which is what we're interested in right now. For future workers, I suggest checking that your drawings are appropriate for the task at hand with other people before wasting time beginning the painstaking process of CAD design in COMSOL.

 

The drawings that I ultimately decided to use are attached below. They detail four layers of stacks, each which arrange 15, 12, 8, and 5 (going from bottom to top) Viton damping springs in an orderly fashion. The numbers have been chosen to keep the load per spring as constant as possible, in order for all springs to oscillate with as close a resonant frequency to each other as possible.

 

Stack_1_Drawing.JPG   Stack_2_Drawing.JPG

 

 Stack_3_Drawing.JPG    Stack_4_Drawing.JPG

 

 Isolation_Stack.JPG

 

After making some minor simplifications, I have finished modeling the top stack:

 

Stack_4_1.png

 

After triangular meshing, before my many failed attempts at mechanical testing:

 

Stack_4_2.png

 

Both the Viton and steel portions have been given their material properties, and so I should be (theoretically ) ready to perform some eigenfrequency calculations on this simplified system. If my predictions are correct, these eigenfrequencies shouldn’t be too far of the eigenfrequencies of the 4-layer stacked system, because of the layout of the springs. I’ve tried doing mechanical tests on the top stack alone, but there hasn’t been much progress yet on this end, mostly because of some boundary value exceptions that COMSOL keeps throwing at me.

 

In the next couple weeks or so, I plan to extend this model to combine all four layers into one completed stack, along with a simple steel base and mirror platform. I also plan to figure out how to make eigenfrequency and transfer function measurements.

 

Lastly, to anyone who is experienced with COMSOL, I am facing two major roadblocks and could really use your help:

 

1) How to import one model into another, merge models together, or copy and paste the same model over and over.

2) Understanding and debugging run-time errors during mechanical testing.

 

If you have any idea how to attack either of these issues, please talk to me! Thanks!

 

ELOG V3.1.3-